OpenStudy (anonymous):

(1/x-3 + 4/x) / (4/x - 1/x-3)

OpenStudy (anonymous):

\[\frac{\frac{1}{x}-3+\frac{4}{x}}{\frac{4}{x}-\frac{1}{x}-3} * \frac{x}{x} = \rightarrow simplify \rightarrow = \frac{3x-5}{3x-3}\]

terenzreignz (terenzreignz):

The trick with these uber complex rational expressions is to reduce the whole thing to an expression with only one fraction bar. Starting on the top: 1/x-3 + 4/x, this is equal to (x + 4x - 12)/x(x - 3) (5x - 12)/x(x - 3), we'll put this to one side... At the bottom: 4/x - 1/x-3, this is equal to (4x - 12 - x)/x(x - 3) (3x - 12)/x(x - 3), this is our denominator. So that entire expression is: [(5x - 12)/x(x-3)] / [(3x - 12)/x(x-3)] This is a bit simpler, but can be further simplified. Since the denominator on top is the same as the denominator on the bottom, they can be cancelled, leaving you with: (5x - 12)/(3x - 12)

OpenStudy (anonymous):

\[ \frac{x}{x} = 1 \], so you can do this anytime

terenzreignz (terenzreignz):

@agentx5 whoops, it appears I have had some difficulty interpreting the smorgasbord of slashes as fraction bars.

OpenStudy (anonymous):

no problem, I was about to go, "err wait a minute" Made me double-check too ;-) ^_^ This is why I keep sounding like a broken record when people don't use parenthesis

OpenStudy (anonymous):

all the slashes are fraction bars

OpenStudy (anonymous):

@zakeyadorsey , is this what you intended to write? \[ \frac{\frac{1}{x}-3+\frac{4}{x}}{\frac{4}{x}-\frac{1}{x}-3} \]

terenzreignz (terenzreignz):

@zakeyadorsey yes, but I don't know what lies under each fraction bar... I'll admit using one lines for uhh complex expressions leads to ambiguities...

OpenStudy (anonymous):

for example: x / 21 + 3x is \[\frac{x}{21}+3x\] whereas: x / ( 21 + 3x ) is \[\frac{x}{21+3x}\]

OpenStudy (anonymous):

Make sense? :-)

OpenStudy (anonymous):

its 1 over x-3 then + 4 over x

terenzreignz (terenzreignz):

Then refer to my first answer :) And feel free to tell me if something isn't clear, or something is wrong, I'm open to your comments

OpenStudy (anonymous):

and that part is the numorator because the denominater is another 4 over x -1 over x-3

terenzreignz (terenzreignz):

but I bet @agentx5 can write it in a clearer manner, I have yet to learn how to write these complex stuff :)

OpenStudy (anonymous):

Then this: \[\frac{\frac{1}{x-3}+\frac{4}{x}}{\frac{4}{x}-\frac{1}{x-3}} \rightarrow simplify \rightarrow \frac{5 x-12}{3 x-12}\]

OpenStudy (anonymous):

Can I haz medals if I teach you all how to do this syntax to make it look like that? ^_^

terenzreignz (terenzreignz):

Sure, if I know how to give medals (me = noob on this site)

OpenStudy (anonymous):

It's two tags

OpenStudy (anonymous):

The first it to make it into an active formula thingy, the LaTeX tags (think HTML tags)

OpenStudy (anonymous):

\[ to open

OpenStudy (anonymous):

\] to close

OpenStudy (anonymous):

Second...

OpenStudy (anonymous):

ok i did not know that

OpenStudy (anonymous):

and yes you can have a medal

OpenStudy (anonymous):

The fraction is created by \frac{numerator}{denominator} It looks like this all together: \[ \frac{numerator}{denominator} \]

OpenStudy (anonymous):

yes i think

terenzreignz (terenzreignz):

I'm listening intently

OpenStudy (anonymous):

Try it! (and if you figure it out a medal would be much appreciated) This also works for polynomial exponents or subtexts: \[\frac{x^{2x+1}}{e^{x^4}} \] \[C_6H_{12}O_6\]

OpenStudy (anonymous):

Without the tags: \frac{x^{2x+1}}{e^{x^4} C_6H_{12}O_6

OpenStudy (anonymous):

ok thanks

terenzreignz (terenzreignz):

now teach me how to give a medal :)

OpenStudy (anonymous):

You put the \[ type tags around the text you want to turn into a formula, basically \frac{}{} is by far the most useful, also sqrt{} (squareroot)

OpenStudy (anonymous):

Clicky the "best answer" button :-D

terenzreignz (terenzreignz):

Thanks :)

OpenStudy (anonymous):

I hope this helps, this LaTeX stuff should really come with a tutorial lol